# Quantum-Limited Acoustoelectric Amplification in a Piezoelectric-2DEG Heterostructure

### v2

@@ I Introduction
Here, we study the quantum limits of amplification for an acoustoelectric phonon amplifier by analyzing the fundamental interactions between the quantum states of a 2D electron gas and piezoelectric acoustic waves. Due to the piezoelectric nature of the acoustic medium, each phonon carries an electric field, which penetrates into the 2DEG and interacts with the electrons there via electric dipole interaction. The electric field thus mediates the electron-phonon coupling Chatterjee et al. (2024). A dc voltage is applied to the 2DEG, effectively pumping the electron distribution and giving rise to an inverted population that amplifies a propagating surface acoustic wave through stimulated emission. The key advantages provided by the 2DEG are an intrinisic conduction-band population, vastly increased mobility, and vastly reduced electric field screening compared to a bulk semiconductor (thus even further enhancing the electron-phonon coupling strength).
@@ II Connecting the Classical and Quantum Pictures
In the 2D regime, near zero temperature and in the absence of the applied dc voltage, the states within the Fermi circle ($k<k_{F}$) will be occupied, while those outside will be unoccupied. If we apply a dc voltage corresponding to an average electron drift momentum $k_{d}$, then the circle shifts in phase space such that it will be centered at $(k_{x},k_{y})=(k_{d},0)$. We thus use a new coordinate in phase space $k_{x}^{\prime}$, defined as $k_{x}^{\prime}=k_{x}-k_{d}$, such that the shifted Fermi circle is centered at $k_{x}^{\prime}=0$. Then, we conceptually deduce that in order for net emission to occur from initial electronic states $k_{x,i}$ to final states $k_{x,f}$, the former must feature an excess of electrons relative to the latter, which is satisfied under the following condition:
@@ III Deriving the Electric Susceptibility
The key step in determining the amplification gain is deriving the first-order electric susceptibility $\chi^{(1)}(-\omega_{0})$, where $\omega_{0}$ is the frequency of the traveling acoustic field. The imaginary part of $\chi^{(1)}$ governs the net phonon emission rate per unit volume per unit intensity of the electric field, while the real and imaginary parts combine with external screening mechanisms to govern the material’s overall dielectric screening, thus inversely controlling the electric field intensity per phonon. In the following subsections, we will analytically derive $\textrm{Im}[\chi^{(1)}(-\omega_{0})]$ in the high-mobility/low-drift-velocity, low-mobility, and high-drift-velocity limits. In addition, we will perform a numerical calculation for the overall range of mobilities and drift velocities. In all regimes, we will start with a model based on the Lindhard theory Lindhard (1970), building on our previous analysis of the susceptibilities in the zero-drift-velocity case Chatterjee et al. (2024). We will then apply the Mermin correction Mermin (1953) to ensure electron number conservation:
@@
where $\chi^{(1)}_{L}(q,\Omega,\gamma)$ represents the Lindhard susceptibility for a given phonon wavevector $q$, effective phonon frequency $\Omega$, and electronic lifetime $\gamma$ (defined as $\gamma=q_{e}/(m\mu)$, where $q_{e}$, $m$, and $\mu$ are the electron charge, effective electron mass, and mobility, respectively). The effective phonon frequency is defined as the phonon frequency in the frame of the drifting electrons, thus yielding $\Omega=(v_{s}-v_{d})q$. Note that in the coming subsections, we will label $\chi^{(1)}(q,-\Omega,\gamma)$ and $\chi^{(1)}_{L}(q,-\Omega,\gamma)$ as $\chi^{(1)}(-\omega_{0})$ and $\chi^{(1)}_{L}(-\omega_{0})$, respectively, for purposes of brevity.
@@ III.1 General Derivation for Imaginary Lindhard Susceptibility
Note that this value is always positive, since $k_{x}^{\prime}>-q/2$ for all $k_{x}^{\prime}\in S^{\prime}$. It is immediately apparent that if $v_{d}<v_{s}$, the detuning is greater for the emissive case, leading the absorptive process to be favored. An example of this is the zero-bias case, where net phonon absorption takes place at any mobility Chatterjee et al. (2024). On the other hand, if $v_{d}>v_{s}$, the the detuning is greater for the absorptive case, and the emissive process is favored. Moreover, the discrepancy between the absorptive and emissive detuning values increases with drift velocity, ensuring a similar increase for the net emission rate with the drift velocity.
@@
The specific methods for deriving the imaginary parts of the Lindhard susceptibility are shown in Appendices A and B (for the high-mobility/low-drift-velocity and low-mobility limits, respectively). The general method is described by the integral in Eq. (LABEL:eq:_low-mobility_chi(1)_integral), which we re-write as follows:
@@
where $f_{\mathrm{net},k_{x}^{\prime}}^{(1)}$ is defined in terms of the average detuning from Eq. (11) and the absorptive and emissive offsets shown in Eqs (8) and (9) as follows:
@@ III.3 High-Mobility, Low-Drift-Velocity Regime
Figure 4: Phase-space diagram (a) of the acoustoelectric amplification process in the high-mobility/low-drift-velocity limit given an acoustic field propagating in the $\hat{x}-$direction with wavevector $q$, along with dimensions of each emissive (red) region (b).
@@ III.4 Low-Mobility Regime
where in the second line, we make the replacement $k_{F}^{2}\rightarrow 2\pi n$ in order to express the imaginary part of $\chi^{(1)}$ as a function of carrier density, in addition to applying the substitution $\gamma=q_{e}/(m\mu)$. The cubic rather than linear dependence on mobility is attributable to the fact that the absorptive and emissive contributions destructively interfere with each other, with the rates of the 2 processes becoming increasingly balanced as the mobility is lowered. The result in terms of carrier density accounts for the high-temperature case $k_{B}T\gg\hbar v_{F}k_{F}$, where the carrier density exceeds the value predicted from the Fermi wavevector due to Fermi-Dirac broadening, demonstrating that $\textrm{Im}[\chi^{(1)}_{L}$] for a given carrier density and mobility is invariant in temperature in both the medium-temperature and high-temperature regimes. This can be attributed to different reasons for the 2 regimes. In the medium-temperature regime, as the thermal energy is increased from $k_{B}T\sim\hbar v_{F}q$ to $k_{B}T\sim\hbar v_{F}k_{F}$, the fraction of electrons participating in electron-phonon interaction increases linearly as $k_{B}T/(\hbar v_{F}q)$ (corresponding to an effective widening of the crescent in Figs. 3), while the average interaction probability for each eligible electron is attenuated by a factor $\hbar v_{F}q/(k_{B}T)$ due to the corresponding attenuation in the Fermi-Dirac occupation number difference between the initial and final electronic states, with the changes cancelling each other out. In the high-temperature regime, increasing the spectral broadening beyond $k_{B}T\sim\hbar v_{F}k_{F}$ causes the average electron-phonon detuning to increase as approximately $\sqrt{k_{B}T/(\hbar v_{F}k_{F})}$, causing the raw interaction probability per electron (i.e., the probability assuming occupied initial and unoccupied final electronic states) to increase by the same factor. However, the average occupation number difference between initial and final electronic states is reduced by the same factor, thus cancelling out the increase in the raw interaction probability.
@@
where we made the replacement $k_{F}^{2}\rightarrow 2\pi n$ in the second step. For the same reason as with the imaginary part of $\chi^{(1)}_{L}$, the variation of the real part of $\chi^{(1)}_{L}$ with carrier density and mobility is identical in the medium-low-mobility ($v_{F}k_{F}\gtrsim\gamma\gtrsim v_{F}q$) and the ultra-low-mobility ($\gamma\gtrsim v_{F}k_{F}$) regimes, particularly because the raw electron-phonon interaction rates for both the real and imaginary cases feature the same variation (i.e., a linear variation) with the electron-phonon detuning (see the discussion in the derivation for the low-mobility result for $\textrm{Im}[\chi^{(1)}]$ for an explanation of how this ensures a mobility-invariant interaction rate per electron per mobility-cubed in the ultra-low-mobility limit).
Figure 6: Numerical results (solid, green) for the real (a) imaginary (b) parts of $\chi^{(1)}(-\omega_{0})$ as functions of the drift velocity $v_{d}$, given a mobility of $1\textrm{ m}^{2}/(\textrm{V}\cdot\textrm{s})$), along with the analytical results for the low-mobility limit $v_{d},v_{F}\ll\gamma/q$ in the low-velocity-mismatch (dotted, red), medium-velocity-mismatch (dashed, blue) and high-velocity-mismatch (dash-dotted, magenta) regimes. Note that the high-mismatch results for both the real and imaginary parts, as well as the medium-mismatch result for the imaginary part, match the corresponding classical results. We assume a phonon angular frequency of $\omega_{0}=2\pi\times 10^{9}\textrm{ s}^{-1}$, speed of sound $v_{s}=4\times 10^{3}$ m/s, a carrier density $n=2\times 10^{15}\textrm{ m}^{-2}$, a carrier effective mass $m=0.067m_{0}$, and a 2DEG thickness $t_{\mathrm{2DEG}}=2\times 10^{-8}\textrm{ m}$.
@@
Figure 6 depicts the real and imaginary parts of the susceptibility in terms of drift velocity in the low-mobility regime. We use the same material parameters as in Fig. 5, except for changing the mobility to $1\textrm{ m}^{2}/(\textrm{V}\cdot\textrm{s})$ for the numerical example. It is interesting to note that our quantum model broadens the Drude model’s divergence for the imaginary part of the susceptibility around $v_{d}=v_{s}$. This provides a more physically complete picture, since a finite carrier density and a finite dipole length (limited by the half-wavelength of the applied ac electric field) ensure that even a maximally polarized electron gas cannot fully cancel out the applied field, thus rendering the divergence of the imaginary part at $v_{d}=v_{s}$ physically untenable.
@@ III.5 High-Drift-Velocity Regime
The real and imaginary parts of the susceptibility in the high-drift-velocity limit are shown in Fig. 7. Note that these are in log-log scale, with the first plot actually depicting the opposite of the real susceptibility (since the real susceptibility in this regime is negative). As expected, both the real and imaginary parts decline with increasing drift velocity in this regime. However, while the real part for any given drift velocity is invariant in mobility, the imaginary part is about 10 times lower for a mobility of 10001000 compared to $100\textrm{ m}^{2}/(\textrm{V}\cdot\textrm{s})$, reflecting the aforementioned inverse variation between the imaginary susceptibility and the mobility.
@@ IV Calculating the Amplifier Gain
Figure 8(a) and (b) show the gain per unit length for the amplifier waveguide for given mobility values in the high-mobility regime, assuming an acoustic field mode depth $L_{z}\sim v_{s}/\omega_{0}$. Counter-intuitively, the peak gain is maximized in the high-drift-velocity regime (at about 5 times the Fermi velocity given our parameters). This is explained by the fact that for $|v_{d}|>v_{F}$, an increasing drift velocity yields a tradeoff: The drop in the imaginary part of $\chi^{(1)}$ reduces the phonon emission rate per unit electric field intensity, while the drop in the dielectric screening increases the field intensity per phonon. In the high-drift-velocity regime, the latter rises faster than the former drops (at rates of $v_{d}^{4}$ and $v_{d}^{-3}$, respectively). As a result, the net effect is for the gain to rise with drift velocity in the high-drift-velocity regime. However, the dielectric screening amplitude eventually reaches its floor once the 2DEG screening roughly cancels out the piezoelectric material’s screening, at which point the electric field intensity per phonon and thus the gain sharply rise. Thereafter, the screening starts to rise again, asymptotically approaching $\epsilon_{p}$, thus causing the gain to decrease from the sharp peak.
@@
Figure 8: Numerical results for the amplifier waveguide gain per unit length $G$ as a function of the drift velocity $v_{d}$ for lower (a) and higher (b) drift velocities, along with amplifier cavity gain per unit length $G_{\mathrm{cav}}$ as a function of the drift speed $|v_{d}|$ for the analogous ranges (c, d), given mobilities (in units of $\textrm{m}^{2}/(\textrm{V}\cdot\textrm{s})$) of 100 (solid, blue), or 1000 (dashed, green). Note that for the cavity in the high-drift-velocity range, the gain is negative (positive) for $v_{d}<297v_{s}$ ($v_{d}>297v_{s}$). We assume a phonon angular frequency of $\omega_{0}=2\pi\times 10^{9}\textrm{ s}^{-1}$, speed of sound $v_{s}=4\times 10^{3}$ m/s, a carrier density $n=2\times 10^{15}\textrm{ m}^{-2}$, a carrier effective mass $m=0.067m_{0}$, a 2DEG thickness $t_{\mathrm{2DEG}}=2\times 10^{-8}\textrm{ m}$, and a mode depth $L_{z}=v_{s}/\omega_{0}$.
@@
It is worth specifically analyzing the dependence of the peak gain on mobility (within the high-mobility range) and carrier density. In the high-drift-velocity regime, the gain is counterintuitively lower for higher mobilities (except in the vicinity of the peak). This is because the imaginary part of the susceptibility, and thus the phonon emission rate per unit electric field intensity, scales inversely with mobility, while the electric field intensity per phonon is roughly invariant in mobility since the dielectric screening is dominated by the mobility-independent real susceptibility. However, the gain peaks at higher values for higher mobilities, due to the fact that the peak corresponds to the zeroing out of the real part of the susceptibility, making the imaginary part dominant in the screening and causing the electric field intensity to scale as the inverse-squared of the imaginary susceptibility. The peak gain thus varies inversely with the imaginary susceptibility, hence increasing with mobility. We can confirm this quantitatively by deriving the optimal drift velocity $v_{d,\mathrm{opt}}$ (i.e., the drift velocity where the real part of the 2DEG screening cancels out the piezoelectric material’s screening):
@@
where we substituted the high-drift-velocity result for the real part of $\chi^{(1)}$ from Eq. (45), along with the relationship $a\sim t_{\mathrm{2DEG}}q$. Note that the optimal drift velocity is invariant in the mobility, matching the plot in Fig. 8(b). On the other hand, the optimal drift velocity increases with the carrier density as $n^{1/2}$, thus making it proportional to the Fermi velocity. The corresponding maximum gain is calculated by substituting this into the imaginary part of $\chi^{(1)}$ in the high-drift-velocity regime from Eq. (46) and then into the gain equation from Eq. (48), yielding:
@@
We now turn to the cavity gain per unit length in the high-mobility regime, depicted in Figs. 8(c) and (d) as a function of the drift speed $|v_{d}|$. This is determined by averaging the forward-propagating and backward-propagating waveguide gains:
@@
Considering the overall range of drift velocities, the low-mobility gain is thus maximized when the 2 denominator terms in Eq. (53) expression match each other, yielding the following optimal drift velocity:

### v3

@@ I Introduction
Here, we study the quantum limits of amplification for an acoustoelectric phonon amplifier by analyzing the fundamental interactions between the quantum states of a 2D electron gas and piezoelectric acoustic waves. Due to the piezoelectric nature of the acoustic medium, each phonon carries an electric field, which penetrates into the 2DEG and interacts with the electrons there via electric dipole interaction. The electric field thus mediates the electron-phonon coupling Chatterjee et al. (2024). A dc voltage is applied to the 2DEG, effectively pumping the electron distribution and giving rise to an inverted population that amplifies a propagating surface acoustic wave through stimulated emission. The key advantages provided by the 2DEG are an intrinisic conduction-band population, vastly increased mobility, and vastly reduced electric field screening compared to a bulk semiconductor (thus even further enhancing the electron-phonon coupling strength). To the last point, it is worth noting that strong acoustoelectric effects have been observed in other low-dimensional materials such as graphene Bandhu et al. (2013); Bandhu and Nash (2016); Hernandez-Minguez et al. (2018); Mou et al. (2024) or transition metal dichalcogenides Preciado et al. (2015); Kalameitsev et al. (2019).
@@ II Connecting the Classical and Quantum Pictures
In the 2D regime, near zero temperature and in the absence of the applied dc voltage, the states within the Fermi circle ($k<k_{F}$) will be occupied, while those outside will be unoccupied. If we apply a dc voltage corresponding to an average electron drift momentum $\bm{k_{d,\mathrm{full}}}=\hat{x}k_{d}+\hat{y}k_{\perp}$ (where $\hat{x}$ is the propagation axis of the acoustic wave and $\hat{y}$ is the transverse in-plane axis), then the circle shifts in phase space such that it will be centered at $(k_{x},k_{y})=(k_{d},k_{\perp})$. We thus use a new coordinate system in phase space $(k_{x}^{\prime},k_{y}^{\prime})$, defined as $k_{x}^{\prime}=k_{x}-k_{d}$ and $k_{y}^{\prime}=k_{y}-k_{\perp}$, such that the shifted Fermi circle is centered at $(k_{x}^{\prime},k_{y}^{\prime})=(0,0)$.
Note that since the phonon travels along the $x$-axis, an electron-phonon interaction conserves the electron’s $y$-momentum while shifting its $x$-momentum. This accords with the intuition that an observer in the frame co-moving with the drifting electrons will not see the transverse component of the wave’s propagation, since the wave’s amplitude is uniform along the transverse axis. Moreover, in a free electron gas, the energy varies separately with the momenta along the two axes (i.e., the variation of the energy with momentum along a given axis is independent of the momentum along the perpendicular axis). Therefore, all functions stemming from the interaction vary only with the $x$-component of the electron momentum (i.e., with $k_{d}$). We thus ignore the $y$-component $k_{\perp}$ in our analysis.
Then, we conceptually deduce that in order for net emission to occur from initial electronic states $k_{x,i}$ to final states $k_{x,f}$, the former must feature an excess of electrons relative to the latter, which is satisfied under the following condition:
@@ III Deriving the Electric Susceptibility
The key step in determining the amplification gain is deriving the first-order electric susceptibility $\chi^{(1)}(-\omega_{0})$, where $\omega_{0}$ is the frequency of the traveling acoustic field. The imaginary part of $\chi^{(1)}$ governs the net phonon emission rate per unit volume per unit intensity of the electric field, while the real and imaginary parts combine with external screening mechanisms to govern the material’s overall dielectric screening, thus inversely controlling the electric field intensity per phonon. In the following subsections, we will analytically derive $\textrm{Im}[\chi^{(1)}(-\omega_{0})]$ in the high-mobility/low-drift-velocity, low-mobility, and high-drift-velocity limits. In addition, we will perform a numerical calculation for the overall range of mobilities and drift velocities. In all regimes, we will start with a model based on the Lindhard theory Lindhard (1970), building on our previous analysis of the susceptibilities in the zero-drift-velocity case Chatterjee et al. (2024). It is worth noting, however, that the Lindhard theory treats scattering-induced electronic state collapse by adding an anti-Hermitian term to the state’s Hamiltonian (corresponding to the state broadening). Although this accounts for the population decay in the initial state, it does not take the impact on the potential final states for the scattered electron into consideration. In the high-mobility limit, the low scattering rate ensures that this impact is negligible. However, for a generic mobility, we will need to apply the Mermin correction Mermin (1953), which ensures electron number conservation by modeling the scattering process through a quantum master equation treatment rather than through an anti-Hermitian effective Hamiltonian:
@@
where $\chi^{(1)}_{L}(q,\Omega,\gamma)$ represents the Lindhard susceptibility for a given phonon wavevector $q$, effective phonon frequency $\Omega$, and electronic decay rate $\gamma$ (defined as $\gamma=q_{e}/(m\mu)$, where $q_{e}$, $m$, and $\mu$ are the electron charge, effective electron mass, and mobility, respectively). The effective phonon frequency is defined as the phonon frequency in the frame of the drifting electrons, thus yielding $\Omega=(v_{s}-v_{d})q$. Note that in the coming subsections, we will label $\chi^{(1)}(q,-\Omega,\gamma)$ and $\chi^{(1)}_{L}(q,-\Omega,\gamma)$ as $\chi^{(1)}(-\omega_{0})$ and $\chi^{(1)}_{L}(-\omega_{0})$, respectively, for purposes of brevity.
@@ III.1 General Derivation for Imaginary Lindhard Susceptibility
Note that this value is always positive, since $k_{x}^{\prime}>-q/2$ for all $k_{x}^{\prime}\in S^{\prime}$. It is immediately apparent that if $v_{d}<v_{s}$, the detuning is greater for the emissive case, leading the absorptive process to be favored. An example of this is the zero-bias case, where net phonon absorption takes place at any mobility Chatterjee et al. (2024). On the other hand, if $v_{d}>v_{s}$, the detuning is greater for the absorptive case, and the emissive process is favored. Moreover, the discrepancy between the absorptive and emissive detuning values increases with drift velocity, ensuring a similar increase for the net emission rate with the drift velocity.
@@
The specific methods for deriving the imaginary parts of the Lindhard susceptibility are shown in Appendices A and B (for the high-mobility/low-drift-velocity and low-mobility limits, respectively). The general method is described by the following integral Chatterjee et al. (2024):
@@
where $\beta$ is a constant function of the 2DEG and piezoelectric material parameters Chatterjee et al. (2024), proportional to the square of the dipole moment for an electron that has undergone interaction with the phonon field and inversely proportional to the 2DEG thickness:
and $f_{\mathrm{net},k_{x}^{\prime}}^{(1)}$ is defined as the net of the inverse complex detuning values for the counter-resonant (absorptive) process for an electron at $k_{x}^{\prime}\in S^{\prime}$ and the corresponding resonant (emissive) process for an electron at $k_{x}\in S$ (see Appendix B). For each electron-phonon interaction process, the inverse complex detuning governs the interaction probability, since it serves as measurement of the closeness of the interaction to resonance. Specifically, the real part corresponds to the probability of a coherent interaction, while the imaginary part corresponds to the probability of a quantum jump (i.e., a phonon absorption or emission event). The imaginary part takes the following form in terms of the average detuning from Eq. (11) and the absorptive and emissive offsets shown in Eqs (8) and (9):
@@ III.2 General Derivation for Real Lindhard Susceptibility
Figure 4: Phase-space diagram (a) of the acoustoelectric amplification process in the high-mobility/low-drift-velocity limit given an acoustic field propagating in the $\hat{x}-$direction with wavevector $q$, along with dimensions of each emissive region (b). Note that in (a), the red-shaded regions represent near-resonance electrons capable of emitting phonons and transitioning to unoccupied states, while the right-hand-side light-blue-shaded region (in between the 2 red regions) represents near-resonance electrons forbidden from emitting phonons since the corresponding final states (represented by the left-hand-side light-blue-shaded region) are occupied. This comports with the fact that the red-shaded regions are subsets of the broader emission-allowed region $S$.
@@ III.4 Low-Mobility Regime
where in the second line, we make the replacement $k_{F}^{2}\rightarrow 2\pi n$ in order to express the imaginary part of $\chi^{(1)}$ as a function of carrier density, in addition to applying the substitution $\gamma=q_{e}/(m\mu)$. The cubic rather than linear dependence on mobility is attributable to the fact that the absorptive and emissive contributions destructively interfere with each other, with the rates of the 2 processes becoming increasingly balanced as the mobility is lowered. The result in terms of carrier density accounts for the high-temperature case $k_{B}T\gg\hbar v_{F}k_{F}$, where the carrier density exceeds the value predicted from the Fermi wavevector due to Fermi-Dirac broadening, demonstrating that $\textrm{Im}[\chi^{(1)}_{L}$] for a given carrier density and mobility is invariant in temperature in both the medium-temperature and high-temperature regimes. This can be attributed to different reasons for the 2 regimes. In the medium-temperature regime, as the thermal energy is increased from $k_{B}T\sim\hbar v_{F}q$ to $k_{B}T\sim\hbar v_{F}k_{F}$, the fraction of electrons participating in electron-phonon interaction increases linearly as $k_{B}T/(\hbar v_{F}q)$ (corresponding to an effective widening of the crescent in Fig. 3), while the average interaction probability for each eligible electron is attenuated by a factor $\hbar v_{F}q/(k_{B}T)$ due to the corresponding attenuation in the Fermi-Dirac occupation number difference between the initial and final electronic states, with the changes cancelling each other out. In the high-temperature regime, increasing the spectral broadening beyond $k_{B}T\sim\hbar v_{F}k_{F}$ causes the average electron-phonon detuning $\left\langle(\Delta\omega)_{k_{x}^{\prime}}\right\rangle$ to increase as approximately $\sqrt{k_{B}T/(\hbar v_{F}k_{F})}$, causing the raw interaction probability per electron (i.e., the probability assuming occupied initial and unoccupied final electronic states) to increase by the same factor. However, the average occupation number difference between initial and final electronic states is reduced by the same factor, thus cancelling out the increase in the raw interaction probability.
@@
where we made the replacement $k_{F}^{2}\rightarrow 2\pi n$ in the second step. For the same reason as with the imaginary part of $\chi^{(1)}_{L}$, the variation of the real part of $\chi^{(1)}_{L}$ with carrier density and mobility is identical in the medium-temperature ($v_{F}k_{F}\gtrsim k_{B}T/\hbar\gtrsim v_{F}q$) and the high-temperature ($k_{B}T/\hbar\gtrsim v_{F}k_{F}$) regimes, particularly because the raw electron-phonon interaction rates for both the real and imaginary cases feature the same variation (i.e., a linear variation) with the average electron-phonon detuning $\left\langle(\Delta\omega)_{k_{x}^{\prime}}\right\rangle$ (see the discussion in the derivation for the low-mobility result for $\textrm{Im}[\chi^{(1)}]$ for an explanation of how this ensures a temperature-invariant interaction rate per electron per mobility-cubed in the high-temperature limit).
Figure 6: Numerical results (solid, green) for the real (a) imaginary (b) parts of $\chi^{(1)}(-\omega_{0})$ as functions of the drift velocity $v_{d}$, given a mobility of $0.3\textrm{ m}^{2}/(\textrm{V}\cdot\textrm{s})$), along with the analytical results for the low-mobility limit $v_{d},v_{F}\ll\gamma/q$ in the low-velocity-mismatch (dotted, red), medium-velocity-mismatch (dashed, blue) and high-velocity-mismatch (dash-dotted, magenta) regimes. Note that the high-mismatch results for both the real and imaginary parts, as well as the medium-mismatch result for the imaginary part, match the corresponding classical results. We assume a phonon angular frequency of $\omega_{0}=2\pi\times 10^{9}\textrm{ s}^{-1}$, speed of sound $v_{s}=4\times 10^{3}$ m/s, a carrier density $n=2\times 10^{15}\textrm{ m}^{-2}$, a carrier effective mass $m=0.067m_{0}$, and a 2DEG thickness $t_{\mathrm{2DEG}}=2\times 10^{-8}\textrm{ m}$.
@@
Figure 6 depicts the real and imaginary parts of the susceptibility in terms of drift velocity in the low-mobility regime. We use the same material parameters as in Fig. 5, except for changing the mobility to $0.3\textrm{ m}^{2}/(\textrm{V}\cdot\textrm{s})$ for the numerical example. It is interesting to note that our quantum model broadens the Drude model’s divergence for the imaginary part of the susceptibility around $v_{d}=v_{s}$. This provides a more physically complete picture, since a finite carrier density and a finite dipole length (limited by the half-wavelength of the applied ac electric field) ensure that even a maximally polarized electron gas cannot fully cancel out the applied field, thus rendering the divergence of the imaginary part at $v_{d}=v_{s}$ physically untenable.
@@ III.5 High-Drift-Velocity Regime
It is also worth considering the medium-temperature and high-temperature cases ($v_{F}k_{F}\gtrsim k_{B}T/\hbar\gtrsim v_{F}q$ and $k_{B}T/\hbar\gtrsim v_{F}k_{F}$, respectively). Although these thermally-broadened regimes are naturally associated with the low-mobility limit (since the thermal broadening sets the baseline for the electronic spectral broadening), the fact that $v_{d}\gg v_{F}$ in the high-drift-velocity limit implies that there may exist a temperature window in the medium-temperature range associated with a high-mobility regime (where the spectral broadening is less than the electron-phonon detuning). It is important to note that as the temperature is varied within this window, the change in the number of transition-allowed electrons is roughly cancelled out by the change in the average Fermi-Dirac occupation number difference between initial and final electronic states, as explained in the previous subsection. On the other hand, the high-temperature range can generally be assumed to fall within the low-mobility regime, since $q\ll k_{F}$, implying that the drift velocities featuring significant gain are close enough to the Fermi velocity such that $v_{d}q\ll v_{F}k_{F}$.
@@
The real and imaginary parts of the susceptibility in the high-drift-velocity limit are shown in Fig. 7. Note that these are in log-log scale, with the first plot actually depicting the opposite of the real susceptibility (since the real susceptibility in this regime is negative). As expected, both the real and imaginary parts decline with increasing drift velocity in this regime. However, while the real part for any given drift velocity is invariant in mobility, the imaginary part is about 10 times lower for a mobility of $1000$ compared to $100\textrm{ m}^{2}/(\textrm{V}\cdot\textrm{s})$, reflecting the aforementioned inverse variation between the imaginary susceptibility and the mobility.
@@ IV Calculating the Amplifier Gain
Figure 8 shows the gain per unit length for the amplifier waveguide for given mobility values in the high-mobility regime, assuming an acoustic field mode depth $L_{z}\sim v_{s}/\omega_{0}$. Counter-intuitively, the peak gain is maximized in the high-drift-velocity regime (at about 5 times the Fermi velocity given our parameters). This is explained by the fact that for $|v_{d}|>v_{F}$, an increasing drift velocity yields a tradeoff: The drop in the imaginary part of $\chi^{(1)}$ reduces the phonon emission rate per unit electric field intensity, while the drop in the dielectric screening increases the field intensity per phonon. In the high-drift-velocity regime, the latter rises faster than the former drops (at rates of $v_{d}^{4}$ and $v_{d}^{-3}$, respectively). As a result, the net effect is for the gain to rise with drift velocity in the high-drift-velocity regime. However, the dielectric screening amplitude eventually reaches its floor once the 2DEG screening roughly cancels out the piezoelectric material’s screening, at which point the electric field intensity per phonon and thus the gain sharply rise. Thereafter, the screening starts to rise again, asymptotically approaching $\epsilon_{p}$, thus causing the gain to decrease from the sharp peak.
@@
Figure 8: Numerical results for the amplifier waveguide gain per unit length $G$ as a function of the drift velocity $v_{d}$ for lower (a) and higher (b) drift velocities, given mobilities (in units of $\textrm{m}^{2}/(\textrm{V}\cdot\textrm{s})$) of 100 (solid, blue), or 1000 (dashed, green). We assume a phonon angular frequency of $\omega_{0}=2\pi\times 10^{9}\textrm{ s}^{-1}$, speed of sound $v_{s}=4\times 10^{3}$ m/s, a carrier density $n=2\times 10^{15}\textrm{ m}^{-2}$, a carrier effective mass $m=0.067m_{0}$, a 2DEG thickness $t_{\mathrm{2DEG}}=2\times 10^{-8}\textrm{ m}$, and a mode depth $L_{z}=v_{s}/\omega_{0}$.
@@
It is worth specifically analyzing the dependence of the peak gain on mobility (within the high-mobility range) and carrier density. In the high-drift-velocity regime, the gain is counterintuitively lower for higher mobilities (except in the vicinity of the peak). This is because the imaginary part of the susceptibility, and thus the phonon emission rate per unit electric field intensity, scales inversely with mobility, while the electric field intensity per phonon is roughly invariant in mobility since the dielectric screening is dominated by the mobility-independent real susceptibility.
Figure 9: Numerical results for the amplifier cavity gain per unit length $G_{\mathrm{cav}}$ as a function of the drift speed $|v_{d}|$ for lower (a) and higher (b) drift speeds, given mobilities (in units of $\textrm{m}^{2}/(\textrm{V}\cdot\textrm{s})$) of 100 (solid, blue), or 1000 (dashed, green). Note that in the high-drift-velocity range, the gain is negative (positive) for $|v_{d}|<297v_{s}$ ($|v_{d}|>297v_{s}$). We assume a phonon angular frequency of $\omega_{0}=2\pi\times 10^{9}\textrm{ s}^{-1}$, speed of sound $v_{s}=4\times 10^{3}$ m/s, a carrier density $n=2\times 10^{15}\textrm{ m}^{-2}$, a carrier effective mass $m=0.067m_{0}$, a 2DEG thickness $t_{\mathrm{2DEG}}=2\times 10^{-8}\textrm{ m}$, and a mode depth $L_{z}=v_{s}/\omega_{0}$.
However, the gain peaks at higher values for higher mobilities, due to the fact that the peak corresponds to the zeroing out of the real part of the susceptibility, making the imaginary part dominant in the screening and causing the electric field intensity to scale as the inverse-squared of the imaginary susceptibility. The peak gain thus varies inversely with the imaginary susceptibility, hence increasing with mobility. We can confirm this quantitatively by deriving the optimal drift velocity $v_{d,\mathrm{opt}}$ (i.e., the drift velocity where the real part of the 2DEG screening cancels out the piezoelectric material’s screening):
@@
where we substituted the high-drift-velocity result for the real part of $\chi^{(1)}$ from Eq. (46), along with the relationship $a\sim t_{\mathrm{2DEG}}q$. Note that the optimal drift velocity is invariant in the mobility, matching the plot in Fig. LABEL:fig:gain(b). On the other hand, the optimal drift velocity increases with the carrier density as $n^{1/2}$, thus making it proportional to the Fermi velocity. The corresponding maximum gain is calculated by substituting this into the imaginary part of $\chi^{(1)}$ in the high-drift-velocity regime from Eq. (47) and then into the gain equation from Eq. (49), yielding:
@@
We now turn to the cavity gain per unit length in the high-mobility regime, depicted in Fig. 9 as a function of the drift speed $|v_{d}|$. This is determined by averaging the forward-propagating and backward-propagating waveguide gains:
@@
Figure 10: Numerical results (solid, green) for the amplifier waveguide gain per unit length $G$ as a function of the drift velocity $v_{d}$, given a mobility of $0.3\textrm{ m}^{2}/(\textrm{V}\cdot\textrm{s})$), along with the analytical results for the low-mobility limit $v_{d},v_{F}\ll\gamma/q$ in the low-velocity-mismatch (dotted, red), medium-velocity-mismatch (dashed, blue) and high-velocity-mismatch (dash-dotted, magenta) regimes. Note that the high-mismatch results match the classical results. We assume a phonon angular frequency of $\omega_{0}=2\pi\times 10^{9}\textrm{ s}^{-1}$, speed of sound $v_{s}=4\times 10^{3}$ m/s, a carrier density $n=2\times 10^{15}\textrm{ m}^{-2}$, a carrier effective mass $m=0.067m_{0}$, and a 2DEG thickness $t_{\mathrm{2DEG}}=2\times 10^{-8}\textrm{ m}$.
@@
Figure 10 depicts the gain per unit length for a mobility of $1\textrm{ m}^{2}/(\textrm{V}\cdot\textrm{s})$ for a broader (a) and zoomed-in (b) range of drift velocities. It is worth noting that the numerically determined gain consistently matches the classical Drude-model results (given by the high-mismatch results). Considering the overall range of drift velocities, the low-mobility gain is maximized when the 2 denominator terms in Eq. (54) expression match each other, yielding the following optimal drift velocity: